# thanks: Corresponding Author

### v1

@@ Abstract
Mass and full width of the heavy scalar molecule $\mathcal{M}$ composed of the mesons $B_{c}^{+}$ and $B_{c}^{-}$ are calculated in QCD sum rule framework. To find its mass and current coupling we apply the two-point sum rule method. The width of the hadronic molecule $\mathcal{M}=B_{c}^{+}B_{c}^{-}$ is evaluated by taking into account its dissociation to $\eta_{c}\eta_{b},J/\psi\Upsilon$, and $B_{c}^{+}B_{c}^{-}$ mesons. Decays into $D$ and $B$ meson pairs with appropriate charges and quantum numbers triggered by $\overline{b}b$ and $\overline{c}c$ annihilations to light quark-antiquarks are also included into analysis. Because the partial widths of $\mathcal{M}$ molecule’s decay channels depend on the strong couplings at $\mathcal{M}$-meson-meson vertices, we estimate them by invoking tools of the three-point sum rule method. Our predictions $m=(12725\pm 85)~{}\mathrm{MeV}$ and $\Gamma[\mathcal{M}]=(155\pm 23)~{}\mathrm{MeV}$ for the parameters of the molecule $\mathcal{M}$ provide useful information for experimental studies of numerous heavy resonances.
@@ I Introduction
Problems of multiquark hadrons, their inner organizations, production and decay mechanisms are among important questions of the quark-parton model and quantum chromodynamic (QCD). Two primary motivations drive the investigation of such structures. Firstly, there is a fundamental theoretical interest in exploring particles that could exist within the quark-gluon model framework. Secondly, and more significantly, these yet hypothetical structures offer a mean to elucidate properties and reconcile them with experimental data. For instance, this concept was utilized in Ref. Jaffe:1976ig to analyze a hierarchy of the masses within the lowest scalar multiplet. According to this assumption, the nonet of the light scalars are four-quark states $q^{2}\overline{q}^{2}$ which accounts for the experimentally observed features of these mesons.
Production of numerous $J^{\mathrm{PC}}=1^{--}$ states by $e^{+}e^{-}$ annihilation in the energy range $3.9-4.4~{}\mathrm{GeV}$ triggered attempts to explain this phenomenon by existence of exotic $c\overline{q}\overline{c}q$ mesons Bander:1975fb . These four-quark particles are made of color-singlet $c\overline{q}$ and $\overline{c}q$ components. Similar concepts regarding the formation of charmonium molecules were also advanced by other researches Voloshin:1976ap ; DeRujula:1976zlg . In this scenario, the four-quark mesons manifest as bound or/and resonant states of the $D$ meson pairs, interacting via conventional light meson exchange.
@@
Over the past two decades, experimental investigations have yielded significant and rich information regarding potential candidates for exotic tetraquark mesons. All such resonances were examined within the frameworks of both diquark-antidiquark and hadronic molecule models. Prominent among recent experimental achievements is the observation of  four $X$ resonances in di-$J/\psi$ and $J/\psi\psi^{\prime}$ mass distributions by the LHCb, ATLAS, and CMS collaborations LHCb:2020bwg ; Bouhova-Thacker:2022vnt ; CMS:2023owd . These resonances are located in the mass interval $6.2-7.3~{}\mathrm{GeV}$ and are presumed to be hidden-charm scalar tetraquarks $cc\overline{c}\overline{c}$. These fully charmed four-quark resonances were also investigated in our previous articles Agaev:2023wua ; Agaev:2023ruu ; Agaev:2023gaq ; Agaev:2023rpj , where we employed the QCD sum rule (SR) method to evaluate their masses and full decay widths. A comparison of our predictions with collected experimental data led to the interpretation of the lightest state, $X(6200)$, as an $\eta_{c}\eta_{c}$ molecule Agaev:2023ruu , while $X(6600)$ is likely a diquark-antidiquark state with axial-vector constituents Agaev:2023wua . A superposition of a diquark-antdiquark state and a hadronic $\chi_{c0}\chi_{c0}$ molecule exhibits parameters consistent with those of the state $X(6900)$ Agaev:2023ruu ; Agaev:2023gaq . The heaviest resonance, $X(7300)$, can be considered as an admixture of a $\chi_{c1}\chi_{c1}$ molecule and the first radial excitation of $X(6600)$ Agaev:2023rpj .
@@
Another significant discovery by LHCb which has profoundly impacted exotic hadron physics, is the axial-vector resonance $T_{\mathrm{cc}}^{+}$, observed in $D^{0}D^{0}\pi^{+}$ mass distribution Aaij:2021vvq ; LHCb:2021auc . With a $cc\overline{u}\overline{d}$ content, it represents the first experimentally discovered open-charm tetraquark candidate. This state has garnered considerable interest within the high energy physics community and has been explored using various theoretical approaches (see, Refs. Agaev:2021vur ; Agaev:2022ast and references therein). In our previous works Agaev:2021vur ; Agaev:2022ast , we evaluated the mass and decay width of this resonance, by modeling it as both a $[cc][\overline{u}\overline{d}]$ diquark-antidiquark and a $D^{0}D^{\ast+}$ molecule states. Our results suggest that $T_{\mathrm{cc}}^{+}$ preferentially possesses a diquark-antidiquark structure, although the computational accuracy does not preclude its interpretation as a hadronic molecule.
@@
These well-established experimental findings have demonstrated that fully heavy tetraquarks are accessible in ongoing experiments, thereby encouraged research efforts dedicated to the thorough analysis of tetraquark mesons composed exclusively of $c$ and $b$ quarks. Exotic structures with hidden charm and bottom, such as $bc\overline{b}\overline{c}$ tetraquarks, form an interesting family of such particles. These states were also investigated intensively in the literature (see, Refs. Agaev:2024wvp ; Agaev:2024mng ; Agaev:2025 and references therein). In our studies Agaev:2024wvp ; Agaev:2024mng ; Agaev:2025 , we examined the scalar, axial-vector, and tensor $bc\overline{b}\overline{c}$ particles within the diquark-antidiquark framework and calculated their masses and full decay widths.
@@
In this article, we model the scalar $bc\overline{b}\overline{c}$ tetraquark as a hadronic molecule $\mathcal{M}=B_{c}^{+}B_{c}^{-}$ and compute its mass and full width. The mass $m$ and current coupling $\Lambda$ of $\mathcal{M}$ are found using the two-point sum rule method Shifman:1978bx ; Shifman:1978by . As will be demonstrated, the hadronic molecule $\mathcal{M}$ is heavier that two-meson thresholds for $\eta_{b}\eta_{c}$, $J/\psi\Upsilon$, and $B_{c}^{+}B_{c}^{-};$consequently, dissociations into these meson pairs represent its kinematically allowed decay channels. Other mechanisms also exist for decays of $\mathcal{M}$ into ordinary mesons Becchi:2020mjz ; Becchi:2020uvq ; Agaev:2023ara . Indeed, owing to $b\overline{b}$ and $c\overline{c}$ annihilations into $\overline{q}q$ and $\overline{s}s$ pairs, $\mathcal{M}$ readily decay into $D^{(\ast)+}D^{(\ast)-}$, $D^{(\ast)0}\overline{D}^{(\ast)0}$, and $D_{s}^{(\ast)+}D_{s}^{(\ast)-}$, as well as $B^{(\ast)+}B^{(\ast)-}$, $B^{(\ast)0}\overline{B}^{(\ast)0}$, and $B_{s}^{(\ast)0}\overline{B}_{s}^{(\ast)0}$ meson pairs. Decay widths for all decay channels of the $\mathcal{M}$ molecule are evaluated within the framework of the three-point sum rule method, which is necessary for determining the strong couplings at the corresponding $\mathcal{M}$-meson-meson vertices.
@@
This work contains the following sections: In Sec. LABEL:sec:Mass, we determine the mass and current coupling of the scalar molecule $\mathcal{M}$. The fall-apart processes $\mathcal{M}\rightarrow\eta_{b}\eta_{c},\ J/\psi\Upsilon$, and $B_{c}^{+}B_{c}^{-}$ are explored in Sec. III. The decays into mesons $D^{(\ast)+}D^{(\ast)-}$, $D^{(\ast)0}\overline{D}^{(\ast)0}$, and $D^{(\ast)+}D^{(\ast)-}$ are considered in Sec. IV. Partial widths of the processes with two $B$ mesons in the final state are calculated in Sec. V. Here, we evaluate the full width of the molecule $\mathcal{M}$ as well. Our conclusions are collected in Sec. VI.
@@ II Spectroscopic parameters of $\mathcal{M}=B_{c}^{+}B_{c}^{-}$
In the hadronic molecule picture the interpolating current for $\mathcal{M}$
@@
The sum rules for $m$ and $\Lambda$ can be found by calculating $\Pi(p)$ using the parameters of $\mathcal{M}$, as well as by employing the quark propagators. The function $\Pi^{\mathrm{Phys}}(p)$ establishes the phenomenological component of SRs, whereas $\Pi^{\mathrm{OPE}}(p)$ forms its QCD side.
The correlator $\Pi^{\mathrm{Phys}}(p)$ is given by the expression
@@
×$\times\mathrm{Tr}\left[\gamma_{5}S_{b}^{bb^{\prime}}(x)\gamma_{5}S_{c}^{b^{\prime}b}(-x)\right],$ (6)
@@
In the case under consideration $\Pi^{\mathrm{OPE}}(p)$ is equal to the invariant amplitude $\Pi^{\mathrm{OPE}}(p^{2})$. To find SRs for the $m$ and $\Lambda$ one equates $\Pi^{\mathrm{OPE}}(p^{2})$ and $\Pi^{\mathrm{Phys}}(p^{2})$, use the Borel transformation and, benefiting from the quark-hadron duality assumption, carry out continuum subtraction Shifman:1978bx ; Shifman:1978by . As a result, one obtains
@@
which are the SRs for the parameters of $\mathcal{M}$. Above $\Pi(M^{2},s_{0})$ is the amplitude $\Pi^{\mathrm{OPE}}(p^{2})$ after the Borel transformation and continuum subtraction, and depend on Borel and continuum subtraction parameters $M^{2}$ and $s_{0}$. In Eq. (7), we used also the short-hand notation $\Pi^{\prime}(M^{2},s_{0})=d\Pi(M^{2},s_{0})/d(-1/M^{2})$.
@@
To perform numerical computations one needs to specify parameters which enter to Eqs. (7) and (8). The masses of the quarks $m_{b}$ and $m_{c}$, and gluon condensate $\langle\alpha_{s}G^{2}/\pi\rangle$ are universal quantities which do not depend on a considering problem:
$m_{b}=4.18_{-0.02}^{+0.03}~{}\mathrm{GeV}\text{,\ }m_{c}=(1.27\pm 0.02)~{}\mathrm{GeV}\text{,}$
Contrary, parameters $M^{2}$ and $s_{0}$ depend on a task under analysis and their choice should meet requirements of the SR calculations. Namely, they have to be fixed in such a way that to ensure $\mathrm{PC}\geq 0.5$, where $\mathrm{PC}$ is the pole contribution defined as
@@
The convergence of the operator product expansion and stability of extracted $m$ and $\Lambda$ on these parameters are also among necessary conditions. In present computations we take into account the dimension $4$ nonperturbative term. Therefore to guarantee convergence of $\mathrm{OPE}$ it is enough to impose a constraint on this term. We restrict this term by the requirement
@@
The spectral parameters of the molecule $\mathcal{M}$ are computed as their average values in domains Eq. (LABEL:eq:Wind1), and amount to
@@
These results coincide with the sum rule predictions at the point $M^{2}=13.5~{}\mathrm{GeV}^{2}$ and $s_{0}=182.5~{}\mathrm{GeV}^{2}$, where $\mathrm{PC}\approx 0.59$. This ensures the prevalence of the pole contribution in the obtained results, and shows ground-state character of the molecule $\mathcal{M}$.

### v2

@@ Abstract
Mass and full width of the heavy scalar molecule $\mathcal{M}$ composed of the mesons $B_{c}^{+}$ and $B_{c}^{-}$ are calculated in the QCD sum rule framework. To find its mass and current coupling, we apply the two-point sum rule method. The width of the hadronic molecule $\mathcal{M}=B_{c}^{+}B_{c}^{-}$ is evaluated by taking into account its dissociation to $\eta_{c}\eta_{b},J/\psi\Upsilon$, and $B_{c}^{+}B_{c}^{-}$ mesons. Decays into $D$ and $B$ meson pairs with appropriate charges and quantum numbers triggered by $\overline{b}b$ and $\overline{c}c$ annihilations to light quark-antiquarks are also included in the analyses. Because the partial widths of $\mathcal{M}$ molecule’s decay channels depend on the strong couplings at $\mathcal{M}$-meson-meson vertices, we estimate them by invoking tools of the three-point sum rule method. Our predictions $m=(12725\pm 85)~\mathrm{MeV}$ and $\Gamma[\mathcal{\ M}]=(155\pm 23)~\mathrm{MeV}$ for the parameters of the molecule $\mathcal{\ M}$ provide useful information for experimental studies of numerous heavy resonances.
@@ I Introduction
Internal organization of multiquark hadrons, their parameters, production and decay mechanisms are among important questions of the quark-parton model and quantum chromodynamic (QCD). Two primary motivations drive the investigation of such structures. Firstly, there is a fundamental theoretical interest in exploring particles that could exist within the quark-gluon model framework. Secondly, and more significantly, these yet hypothetical structures offer a mean to elucidate properties of the particles and reconcile them with experimental data. For instance, this concept was utilized in Ref. Jaffe:1976ig to analyze a mass hierarchy within the lowest scalar multiplet. According to this work, the nonet of the light scalars are four-quark states $q^{2}\overline{q}^{2}$ which accounts for the experimentally observed features of these mesons.
Production of numerous $J^{\mathrm{PC}}=1^{--}$ states by $e^{+}e^{-}$ annihilation in the energy range $3.9-4.4~\mathrm{GeV}$ triggered attempts to explain this phenomenon by existence of exotic $c\overline{q}\overline{c}q$ mesons Bander:1975fb . These four-quark particles are made of color-singlet $c\overline{q}$ and $\overline{c}q$ components. Similar concepts regarding the formation of charmonium molecules were also advanced by other researchers Voloshin:1976ap ; DeRujula:1976zlg . In this scenario, the four-quark mesons manifest as bound or/and resonant states of the $D$ meson pairs, interacting via conventional light meson exchange.
@@
Over the past two decades, experimental investigations collected significant and rich information regarding potential candidates to exotic tetraquark mesons. All such resonances were examined within the frameworks of both diquark-antidiquark and hadronic molecule models. Prominent among recent experimental achievements is the discovery of  four $X$ resonances in di-$J/\psi$ and $J/\psi\psi^{\prime}$ mass distributions by the LHCb, ATLAS, and CMS collaborations LHCb:2020bwg ; Bouhova-Thacker:2022vnt ; CMS:2023owd . These resonances are located in the mass interval $6.2-7.3~\mathrm{GeV}$ and are presumed to be hidden-charm scalar tetraquarks $cc\overline{c}\overline{c}$. These fully charmed four-quark resonances were also investigated in our previous articles Agaev:2023wua ; Agaev:2023ruu ; Agaev:2023gaq ; Agaev:2023rpj , where we employed the QCD sum rule (SR) method to evaluate their masses and full decay widths. A comparison of our predictions with available experimental data led to the interpretation of the lightest state, $X(6200)$, as an $\eta_{c}\eta_{c}$ molecule Agaev:2023ruu , while $X(6600)$ is likely a diquark-antidiquark state with axial-vector constituents Agaev:2023wua . A superposition of a diquark-antdiquark state and a hadronic $\chi_{c0}\chi_{c0}$ molecule exhibits parameters consistent with those of the state $X(6900)$ Agaev:2023ruu ; Agaev:2023gaq . The heaviest resonance, $X(7300)$, can be considered as an admixture of a $\chi_{c1}\chi_{c1}$ molecule and the first radial excitation of $X(6600)$ Agaev:2023rpj .
@@
Another significant discovery by LHCb which has profoundly impacted exotic hadron physics, is the axial-vector resonance $T_{\mathrm{cc}}^{+}$, observed in $D^{0}D^{0}\pi^{+}$ mass distribution Aaij:2021vvq ; LHCb:2021auc . With a $cc\overline{u}\overline{d}$ content, it represents the first experimentally discovered open-charm tetraquark candidate. This state has garnered considerable interest within the high energy physics community and has been explored using various theoretical approaches (see, Refs. Agaev:2021vur ; Agaev:2022ast and references therein). In our previous works Agaev:2021vur ; Agaev:2022ast , we evaluated the mass and decay width of this resonance, by modeling it as both a $cc][\overline{u}\overline{d}]$ diquark-antidiquark and a $D^{0}D^{\ast+}$ molecule state. Our results suggest that $T_{\mathrm{cc}}^{+}$ preferentially possesses a diquark-antidiquark structure, although the computational accuracy does not preclude its interpretation as a hadronic molecule.
@@
These well-established experimental findings have demonstrated that fully heavy tetraquarks are accessible in ongoing experiments, thereby encouraged research efforts dedicated to the thorough analysis of tetraquark mesons composed exclusively of $c$ and $b$ quarks. Exotic structures with hidden charm and bottom, such as $bc\overline{b}\overline{c}$ tetraquarks, form an interesting family of such particles. These states were investigated intensively in the literature (see, Refs. Agaev:2024wvp ; Agaev:2024mng ; Agaev:2025 and references therein). In our studies Agaev:2024wvp ; Agaev:2024mng ; Agaev:2025 , we examined the scalar, axial-vector, and tensor $bc\overline{b}\overline{c}$ particles within the diquark-antidiquark framework and calculated their masses and full decay widths.
@@
In this article, we model the scalar $bc\overline{b}\overline{c}$ tetraquark as a hadronic molecule $\mathcal{M}=B_{c}^{+}B_{c}^{-}$ and compute its parameters. The mass $m$ and current coupling $\Lambda$ of $\mathcal{M}$ are found using the two-point sum rule method Shifman:1978bx ; Shifman:1978by . As it will be demonstrated, the hadronic molecule $\mathcal{M}$ is heavier than two-meson thresholds $\eta_{b}\eta_{c}$, $J/\psi\Upsilon$, and $B_{c}^{+}B_{c}^{-}$; consequently, dissociations into these meson pairs represent its kinematically allowed decay channels. Other mechanisms also exist for decays of $\mathcal{M}$ into ordinary mesons Becchi:2020mjz ; Becchi:2020uvq ; Agaev:2023ara . Indeed, owing to $b\overline{b}$ and $c\overline{c}$ annihilations into $\overline{q}q$ and $\overline{s}s$ pairs, $\mathcal{M}$ readily decays into $D^{(\ast)+}D^{(\ast)-}$, $D^{(\ast)0}\overline{D}^{(\ast)0}$, and $D_{s}^{(\ast)+}D_{s}^{(\ast)-}$, as well as $B^{(\ast)+}B^{(\ast)-}$, $B^{(\ast)0}\overline{B}^{(\ast)0}$, and $B_{s}^{(\ast)0}\overline{B}_{s}^{(\ast)0}$ meson pairs. Decay widths for all decay channels of the $\mathcal{M}$ molecule are evaluated within the framework of the three-point sum rule method, which is necessary for determining the strong couplings at the corresponding $\mathcal{M}$-meson-meson vertices.
@@
This work contains the following sections: In Sec. II, we determine the mass and current coupling of the scalar molecule $\mathcal{M}$. The fall-apart processes $\mathcal{M}\rightarrow\eta_{b}\eta_{c},\ J/\psi\Upsilon$, and $B_{c}^{+}B_{c}^{-}$ are explored in Sec. III. The decays into mesons $D^{(\ast)+}D^{(\ast)-}$, $D^{(\ast)0}\overline{D}^{(\ast)0}$, and $D^{(\ast)+}D^{(\ast)-}$ are considered in Sec. IV. Partial widths of the processes with two final-state $B$ mesons are calculated in Sec. V. Here, we evaluate the full decay width of the molecule $\mathcal{M}$ as well. Our conclusions are collected in Sec. VI.
@@ II Spectroscopic parameters of $\mathcal{M}=B_{c}^{+}B_{c}^{-}$
In the hadronic molecule picture the interpolating current for $\mathcal{M}$ is
@@
The sum rules for $m$ and $\Lambda$ can be found by calculating $\Pi(p)$ using the physical parameters of $\mathcal{M}$ which establishes the phenomenological component of SRs $\Pi^{\mathrm{Phys}}(p)$. The correlator $\Pi^{\mathrm{Phys}}(p)$ is given by the expression
@@
$\times\mathrm{Tr}\left[\gamma_{5}S_{b}^{bb^{\prime}}(x)\gamma_{5}S_{c}^{b^{\prime}b}(-x)\right],$ (6)
@@
In the case under consideration $\Pi^{\mathrm{OPE}}(p)$ is equal to the invariant amplitude $\Pi^{\mathrm{OPE}}(p^{2})$. To find SRs for the $m$ and $\Lambda$ one equates $\Pi^{\mathrm{OPE}}(p^{2})$ and $\Pi^{\mathrm{Phys}}(p^{2})$, uses the Borel transformation and, benefiting from the quark-hadron duality assumption, carries out continuum subtraction Shifman:1978bx ; Shifman:1978by . As a result, one obtains
@@
which are the SRs for the parameters of $\mathcal{M}$. Above $\Pi(M^{2},s_{0})$ is the amplitude $\Pi^{\mathrm{OPE}}(p^{2})$ after the Borel transformation and continuum subtraction, and it depends on Borel and continuum subtraction parameters $M^{2}$ and $s_{0}$.
@@
$\Pi(M^{2},s_{0})$ $=$ $\int_{4\mathcal{(}m_{b}+m_{c})^{2}}^{s_{0}}ds\rho(s)e^{-s/M^{2}}+\Pi(M^{2}).$
To perform numerical computations, one needs to specify parameters which enter to Eqs. (7) and (8). The masses of the quarks $m_{b}$ and $m_{c}$, and gluon condensate $\langle\alpha_{s}G^{2}/\pi\rangle$ are universal quantities which do not depend on a considering problem:
$m_{b}=(4.183\pm 0.007)\ \mathrm{GeV}\text{,\ }m_{c}=(1.2730\pm 0.0046)~\mathrm{GeV}\text{,}$
@@
In contrast, parameters $M^{2}$ and $s_{0}$ depend on a task under analysis and their choice should meet requirements of the SR calculations. Namely, they have to be fixed in such a way that to ensure $\mathrm{PC}\geq 0.5$, where $\mathrm{PC}$ is the pole contribution defined as
@@
The convergence of the operator product expansion and stability of extracted quantities $m$ and $\Lambda$ on these parameters are also among necessary conditions. In present computations we take into account the dimension-$4$ nonperturbative term. Therefore to guarantee convergence of $\mathrm{OPE}$ it is enough to impose a constraint on this term. We restrict this term by the requirement
@@
The spectral parameters of the molecule $\mathcal{M}$ are computed as their average values in the domains Eq. (LABEL:eq:Wind1), and amount to
@@
These results coincide with the sum rule predictions at the point $M^{2}=13.5~\mathrm{GeV}^{2}$ and $s_{0}=182.5~\mathrm{GeV}^{2}$, where the pole term forms $\mathrm{PC}\approx 59\%$ of the whole contribution. This ensures the prevalence of the pole contribution in the obtained results, and shows ground-state character of the molecule $\mathcal{M}$ in its class. As the function of the parameters $M^{2}$ and $s_{0}$ the mass of the molecule $\mathcal{M}$ is depicted in Fig. 2.